OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store?

OpenStudy (anonymous):

Let the price of an individual CD be "x" and the price of an individual DVD be "y" You get the following two equations: 3x + 5y = 137 2x + 7y = 161 If you have two equations, you can solve for two variables. Solve for x in the first equation: x = (137-5y)/3 Then substitute this in for "x" in the 2nd equation. 2[ (137-5y) / 3 ] + 7y = 161 Now you can solve for "y" (274/3) - (10/3)*y + 7y = 161 (11/3)*y = 161 - (274/3) y = 19 Now that you have "y", you can solve for "x" 2x + 7(19) = 161 x = 14 CDs cost $14 and DVDs cost $19 each